1. If the N2 amount is entirely used to produce ammonia gas, then how many kg’s of H2 we require and how many Kg’s of ammonia gas will be produced?



2. If we need to produce calcium carbonate from that release carbon dioxide gas, how many kg’s of calcium oxide we need to purchase from market?



3. Critical thinking Question:



If we want to produce urea from the above mentioned system, how much Urea we can theoretically produce?




1
Expert's answer
2022-01-10T17:01:04-0500

N2 + 3H2 = 2NH3


28g + 6g = 34g


28ton + 6ton = 34 ton


28/34ton = 0.824 ton N2 + 6/34 = 0.176 ton H2 will produce 34/34 = 1 ton NH3


0.824 ton N2


0.176 ton H2